(a) If an insect with a radius of 0.289m accelerates uniformly from rest to an angular speed of 78.0 rev/min in 2.00 sec, what is its tangential acceleration? (b) What is the tangential velocity of the insect when the disk reaches its maximum speed? What are the insect's (c) tangential acceleration (d) centripetal acceleration (e) total acceleration one second after it starts from rest?


1
Expert's answer
2021-08-26T19:04:22-0400

a)


"a_t=\frac{78(2\pi)}{2(60)}0.289=1.18\frac{m}{s^2}"

b)


"v_t=\frac{78(2\pi)}{(60)}0.289=2.36\frac{m}{s}"

c)


"a_t=1.18\frac{m}{s^2}"

d)


"v_t=\frac{78(2\pi)}{2(60)}0.289=1.18\frac{m}{s}\\
a_c=\frac{1.18^2}{0.289}=4.82\frac{m}{s^2}"

e)


"a=\sqrt{1.18^2+4.82^2}=4.96\frac{m}{s^2}"